OpenStudy (anonymous):

How many 3-letter codes are possible using the first 7 letters of the alphabet if letters can be repeated?

OpenStudy (anonymous):

You have 3 "slots" into which you can put letters. In each slot you can put 7 letters. Therefore there are: \[7 * 7 * 7 = 7^3 = 343 \textrm{ 3-letter codes}\]